# Fixed-sparsity matrix approximation from matrix-vector products

### v1

@@ 3.2 Proof of Theorem 2
$\leq\operatorname{\mathbb{P}}\Bigl{[}\big{|}\mathbf{x}_{i}^{\mathsf{T}}\mathbf{x}_{j}-[\mathbf{T}]_{i,j}\big{|}<\alpha\|\mathbf{v}_{i}\|_{2}\|\mathbf{v}_{j}\|_{2}\sqrt{k}\Bigr{]}<\alpha.$
Since $\alpha=1/100$ we have Equation 9, which proves Equation 8 under the assumptions in Equation 7.
@@
We now relax the assumption $c_{5}/\varepsilon\in\mathbb{Z}$. For arbitrary $\varepsilon>0$, define $\varepsilon^{\prime}=c_{5}/\lceil c_{5}/\varepsilon\rceil$. Then $c_{5}/\varepsilon^{\prime}\in\mathbb{Z}$ and $\varepsilon^{\prime}\leq\varepsilon$. Moreover, if $\varepsilon\leq c_{5}/2$, then $\varepsilon\leq c_{5}\varepsilon^{\prime}/(c_{5}-\varepsilon^{\prime})\leq 2\varepsilon^{\prime}$. We have set $c=\min\{1/c_{6},c_{5}/2\}$. Therefore, since $1/\varepsilon\leq 1/\varepsilon^{\prime}$, if
@@
then the conditions in Equation 7 hold (with $\varepsilon^{\prime}$), and so we have the result in Equation 8 (with $\varepsilon^{\prime}$). Since $\varepsilon\leq 2\varepsilon^{\prime}$, this implies there is a distribution on matrices and sparsity patterns for which the result of the theorem holds.
@@ 4 Comparison with coloring methods
A number of methods for sparse-matrix and operator recovery based on graph colorings have been proposed [CPR74, CM83, SLO13, SO21, FSS21, etc.]. [^4] These methods are sometimes called probing methods in the matrix function trace estimation literature. To the best of our knowledge, such methods were first considered in the 1970s, and are based on the following observation: Partition column indices $[d]$ into sets $\{C_{1},\ldots,C_{k}\}$ such that the columns of $\mathbf{A}$ corresponding to any given $C_{i}$ have disjoint support. Then we can recover all of the columns in a given set $C_{i}$ with a single matrix-vector product: a vector which is supported only on $C_{i}$. The sets $C_{i}$ can be obtained by coloring a graph. In particular, form a graph on $d$ vertices, where there is an edge between vertices $i$ and $j$ if and only if the $i$-th and $j$-th columns of $\mathbf{A}$ have overlapping support. A $k$-coloring of this graph gives the partition $\{C_{1},\ldots,C_{k}\}$.
@@ 4.1 Analysis of coloring methods on banded matrices
For any integer $b\geq 0$, let $\mathbf{S}\in\{0,1\}^{d\times d}$ be the sparsity pattern of banded matrices of bandwidth $s=2b+1$; that is, $[\mathbf{S}]_{i,j}=\mathds{1}(|i-j|\leq b)$. For convenience, we will assume $d=ks$, for some integer $k\geq 1$. This sparsity pattern yields a natural coloring-based partitioning of $[d]$ is into the sets $C_{i}=\{i+js:j\in[k]\}$ for $i\in[s]$.
@@
Matrix-vector products with $\mathbf{A}^{(i)}$ can be computed with a single product to $\mathbf{A}$. Thus, using $s$ matrix-vector products, we obtain an unbiased estimator for $\mathbf{S}\circ\mathbf{A}$ with expected squared error $\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}$. Averaging $t$ independent copies of this estimator will reduce the variance by a factor of $t$. Hence, using $m\geq s/\varepsilon$ matrix-vector products, one obtains an algorithm with expected squared error bounded by $\varepsilon\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}$. While this algorithm is well-known in the literature [SLO13, FSS21, etc.], to the best of our knowledge, an analysis like the one described here has not been written down. As we discuss in the next section, this is marginally better than Theorem 1.
@@ 5 Numerical Experiments
In this section we provide several numerical experiments which illustrate the performance of Algorithm 1. These problems are modeled after similar problems from the literature. Code to reproduce the figures can be found at https://github.com/tchen-research/fixed_sparsity_matrix_recovery.
@@ 5.1 Model problem
We consider the matrix $\mathbf{A}=\mathbf{M}^{-1}$, where $\mathbf{M}=\operatorname{tridiag}(-1,4,-1)$. This class of matrices exhibits exponential decay away from the diagonal and was used in experiments in past work [BS15, FSS21].
@@ 5.2 Trefethen Primes
We let $\mathbf{A}=\mathbf{M}^{-1}$, where $\mathbf{M}$ be the $1000\times 1000$ matrix whose entries are zero everywhere except for the primes $2,3,5,7,\ldots,7919$ along the main diagonal and the number 1 in all the positions $[\mathbf{B}]_{i,j}$ with $|i–j|\in\{1,2,4,8,\ldots,512\}$. [^5] In problem 7 of the “A Hundred-dollar, Hundred-digit Challenge” in SIAM News, readers are asked to compute the (1,1) entry of a larger but analogously defined matrix $\mathbf{A}$ to 100 digits of accuracy [Tre02]. An example (somewhat different from our example) involving this matrix was used in [PN23].
@@
For $b>0$, we defined a sparsity pattern $\mathbf{S}$ to be such that, for each $t\in\{1,2,4,8,\ldots,512\}$, $[\mathbf{S}]_{i,j}=1$ whenever $|i-j\pm t|\leq b$ and zero otherwise. In other words, the sparsity pattern consist of bandwidth $12b+$ 1 bands centered nonzero entries of $\mathbf{M}=\mathbf{A}^{-1}$. This is illustrated in Figure 3.
The results are shown in Figure 4. Here, the convergence of Algorithm 1 is somewhat better than the upper bound Theorem 1. This is because many rows of the sparsity pattern have far fewer than $s$ entries. From the proof of Theorem 1, it is clear how to obtain an exact characterizaton of the epxected squared error.
@@ 6 Outlook
It is clear that there is potential for combining coloring methods with algorithms such as Algorithm 1 in order to avoid paying for large off-sparsity entries in parts of a matrix while simultaneously maintaining the robustness to noise enjoyed by Algorithm 1. One approach to combing these two paradigms is to use adaptive queries to identify portions of the matrix with large-mass, and then find a coloring based on this information. We believe further study in this direction may yield algorithms which work better in many practical situations. Of course, our lower bound Theorem 2 shows that only constant factors can be improved for some families of problem instances.
The present paper focuses only on the Frobenius norm. It would be valubale to understand the analogous problem in other norms such as the matrix 2-norm. For other norms, $\mathbf{S}\circ\mathbf{A}$ is not necessarily best approximation to $\mathbf{A}$ with sparsity $\mathbf{S}$. For instance, if $\mathbf{A}=[1,1;1,1]$ and $\mathbf{S}=[1,0;0,0]$, then
@@
Loosely speaking, minimizing the operator-norm approximation error requires the rows of the error matrix to be small and un-aligned, whereas the Frobenius norm problem only requires them to be small.
@@ Acknowledgements
Appendix A Lower Bound Lemmas
@@
A.1 Adaptive queries to Wishart matrices
In this section we provide a proof of Proposition 3, which is essentially [BHSW20, Lemma 3.4]. This is mostly included for completeness. First, however, we consider what happens after a single non-adaptive query.
@@ A.1 Adaptive queries to Wishart matrices
Suppose $\mathbf{G}\sim\operatorname{Gaussian}(d)$. Let $\mathbf{x}$ be a unit-length query chosen independently of $\mathbf{G}$, and define $\mathbf{y}=\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{x}$. Then, there is is an $n\times n$ orthonormal matrix $\hat{\mathbf{X}}=[\mathbf{x}~{}\mathbf{X}]$, constructed solely as a function of $\mathbf{x}$, and a matrix $\mathbf{H}\sim\operatorname{Gaussian}(n-1)$ independent of $\mathbf{x}$ and $\mathbf{y}$ such that such that
@@
The first row and column of this matrix depend only on $\mathbf{x}$ and $\mathbf{y}$. We will now show $\mathbf{X}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{X}$ consists of a matrix depending on $\mathbf{x}$ and $\mathbf{y}$ and Gaussian matrix independent of $\mathbf{x}$ and $\mathbf{y}$.
@@
$\hat{\mathbf{X}}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{G}\hat{\mathbf{X}}=\begin{bmatrix}\mathbf{x}^{\mathsf{T}}\mathbf{y}&\mathbf{y}^{\mathsf{T}}\mathbf{X}\\ \mathbf{X}^{\mathsf{T}}\mathbf{y}&(\mathbf{x}^{\mathsf{T}}\mathbf{y})^{-2}\mathbf{X}^{\mathsf{T}}\mathbf{y}\mathbf{y}^{\mathsf{T}}\mathbf{X}\end{bmatrix}+\begin{bmatrix}0&\mathbf{0}_{1,n-1}\\ \mathbf{0}_{n-1,1}&\mathbf{X}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{R}\mathbf{R}^{\mathsf{T}}\mathbf{G}\mathbf{X}\end{bmatrix}.$
It remains to show $\mathbf{R}^{\mathsf{T}}\mathbf{G}\mathbf{X}$ is a $(n-1)\times(n-1)$ Gaussian matrix independent of $\mathbf{x}$ and $\mathbf{y}$.
First, note that since ${\mathbf{X}}$ is chosen only based on $\mathbf{x}$ (which is independent of $\mathbf{G}$), $\mathbf{G}\hat{\mathbf{X}}$ consists of iid Gaussians. Thus, the columns of $\mathbf{G}{\mathbf{X}}$ are mutually independent of one another and $\mathbf{x}$, and hence $\mathbf{G}\mathbf{X}$ is mutually independent of $\mathbf{x}$ and $\mathbf{y}$. Finally, since $\mathbf{R}$ depends only on $\mathbf{r}=(\mathbf{x}^{\mathsf{T}}\mathbf{y})^{-1}\mathbf{G}\mathbf{x}$, $\mathbf{R}$ is independent of $\mathbf{G}\mathbf{X}$. Thus, $\mathbf{R}^{\mathsf{T}}\mathbf{G}\mathbf{X}$ has iid Gaussian entries independent of $\mathbf{x}$ and $\mathbf{y}$ (and hence any matrices constructed solely from $\mathbf{x}$ and $\mathbf{y}$). ∎
@@
We proceed by induction. Suppose, at step $t$ the result of the lemma holds. Let $\mathbf{x}_{t+1}$ be a query chosen based solely on $\mathbf{x}_{1},\ldots,\mathbf{x}_{t}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t}$ and hence independent of $\mathbf{G}_{t}$.
Then,
@@
Let $\mathbf{x}$ denote the bottom $d-t$ entries of $\mathbf{V}^{\mathsf{T}}\mathbf{x}_{t+1}$, normalized to have length 1. By Lemma 3 the query to $\mathbf{G}_{t}^{\mathsf{T}}\mathbf{G}_{t}$ results in a factorization
@@
where $\mathbf{\Delta}$ and $\hat{\mathbf{X}}$ are constructed solely as functions of $\mathbf{x}$ and $\mathbf{G}_{t}^{\mathsf{T}}\mathbf{G}_{t}\mathbf{x}$ (and hence of $\mathbf{x}_{1},\ldots,\mathbf{x}_{t+1}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t+1}$), $\hat{\mathbf{X}}$ is orthonormal, and $\mathbf{H}\sim\operatorname{Gaussian}(d-t-1)$ is independent of $\mathbf{x}$ (and hence of $\mathbf{x}_{1},\ldots,\mathbf{x}_{t+1}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t+1}$).
@@
A.2 Anti-concentration for sums

### v3

@@ 3.2 Proof of Theorem 2
$\leq\operatorname{\mathbb{P}}\Bigl{[}\big{|}\mathbf{x}_{i}^{\mathsf{T}}\mathbf{x}_{j}-[\mathbf{T}]_{i,j}\big{|}<(\alpha/2)\|\mathbf{v}_{i}\|_{2}\|\mathbf{v}_{j}\|_{2}\sqrt{r-m}\Bigr{]}<\frac{\alpha}{2}.$
Now, by Markov’s inequality, we have that
$\operatorname{\mathbb{P}}\Biggl{[}\sum_{(i,j)\in M}\mathds{1}[(i,j)\not\in P]\geq\frac{1}{\alpha}\cdot\frac{\alpha}{2}|M|\Biggr{]}\leq\alpha.$
Since $|M|\geq 2c_{1}ds$, this then implies that
$\operatorname{\mathbb{P}}\bigl{[}E\bigr{]}=\operatorname{\mathbb{P}}\biggl{[}|P|\geq c_{1}ds\biggr{]}\geq\operatorname{\mathbb{P}}\biggl{[}|P|\geq\frac{|M|}{2}\biggr{]}\geq 1-\alpha.$
This proves the result. ∎
With Lemma 3, the proof if Theorem 2 is straightforward. The basic idea is that if $r=d$ and $\alpha$ is constant then $\|\mathbf{S}\circ\mathbf{A}-\widetilde{\mathbf{A}}\|_{\mathsf{F}}^{2}$ is typically $\Omega(sd^{2})$. However, by simple computation, it is not hard to see that $\|\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\leq\|\mathbf{A}\|_{\mathsf{F}}^{2}$ is typically $O(d^{3})$. Thus, if we set $d=O(s/\varepsilon)$, we typically will have that $\|\mathbf{S}\circ\mathbf{A}-\widetilde{\mathbf{A}}\|_{\mathsf{F}}^{2}>\varepsilon\|\mathbf{S}\circ\mathbf{A}\|$ as long as $m\leq c_{1}d=O(s/\varepsilon)$.
Proof of Theorem 2.
Fix $\gamma\in(0,1)$ and let $c_{1},c_{2},c_{3}$ be the constants from Lemma 3. We will make the following assignments: [^3] We have labeled the constants so that if $c_{i}$ depends on $c_{j}$, then $i>j$.
$\alpha$ $=\frac{1}{25},$ $b_{1}$ $=50,$ $b_{2}$ $=\frac{c_{2}\alpha^{2}(1-c_{1})}{6b_{1}},$ $b_{3}$ $=\max\left\{2,\frac{(1-c_{1})b_{2}\alpha^{2}}{c_{3}}\right\},$ $C_{1}$ $=c_{1}b_{2}.$
Fix $\varepsilon>0$. We will show that if
$m\leq C_{1}\frac{s}{\varepsilon},\qquad\frac{b_{2}}{\varepsilon}\in\mathbb{Z},\qquad\varepsilon<\frac{1}{b_{3}}\leq\frac{1}{2},$ (7)
The assumption $b_{2}/\varepsilon\in\mathbb{Z}$ will subsequently be removed.
Towards this end, assume Equation 7, and set
$d=b_{2}\frac{s}{\varepsilon},\quad r=d,\quad\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G},\quad\mathbf{G}\sim\operatorname{Gaussian}(r,d).$
Define events,
$E=\Big{\{}\|\mathbf{S}\circ\mathbf{A}-\widetilde{\mathbf{A}}\|_{\mathsf{F}}^{2}\geq c_{2}\alpha^{2}ds(r-m)\Big{\}},\qquad F=\Big{\{}\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\leq b_{1}d^{3}\Big{\}}.$
By assumption, $m\leq C_{1}s/\varepsilon=c_{1}d$ and
$r-m\geq(1-c_{1})d=(1-c_{1})\cdot b_{2}\frac{s}{\varepsilon}>(1-c_{1})b_{2}\frac{1}{b_{3}}=\frac{c_{3}}{\alpha^{2}}.$
Therefore, applying Lemma 3, we have that $\operatorname{\mathbb{P}}[E]\geq 49/50$.
It is easy to show that $\operatorname{\mathbb{E}}[\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}]\leq d^{3}$ (see Fact 2 for a derivation), so since $b_{1}=50$, an application of Markov’s inequality implies $\operatorname{\mathbb{P}}[F]\geq 49/50$.
Note that
$6\varepsilon b_{1}d^{3}\leq 6\varepsilon b_{1}d\cdot b_{2}\frac{s}{\varepsilon}\cdot\frac{r-m}{1-c_{1}}\leq c_{2}\alpha^{2}ds(r-m).$
By a union bound, both $E$ and $F$ hold with probability at least $24/25$. Then since $\mathbf{S}\circ\mathbf{A}$ and $\widetilde{\mathbf{A}}$ have disjoint support and $\varepsilon<1/2$, as noted in Equation 1,
$\|\mathbf{S}\circ\mathbf{A}-\widetilde{\mathbf{A}}\|_{\mathsf{F}}^{2}\geq(1+6\varepsilon)\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\geq(1+2\varepsilon)^{2}\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}.$
@@
We now relax the assumption $c_{5}/\varepsilon\in\mathbb{Z}$. For arbitrary $\varepsilon>0$, define $\tilde{\varepsilon}=b_{2}/\lceil b_{2}/\varepsilon\rceil$. Then $b_{2}/\tilde{\varepsilon}\in\mathbb{Z}$ and $\tilde{\varepsilon}\leq\varepsilon$. Moreover, if $\varepsilon\leq b_{2}/2$, then $\varepsilon\leq b_{2}\tilde{\varepsilon}/(b_{2}-\tilde{\varepsilon})\leq 2\tilde{\varepsilon}$. Set $c=\min\{1/b_{3},b_{2}/2\}$. Therefore, since $1/\varepsilon\leq 1/\tilde{\varepsilon}$, if
@@
then the conditions in Equation 7 hold (with $\tilde{\varepsilon}$), and so we have the result in Equation 8 (with $\tilde{\varepsilon}$). Since $\varepsilon\leq 2\tilde{\varepsilon}$, this implies there is a distribution on matrices and sparsity patterns for which the result of the theorem holds.
@@ 4 Comparison with coloring methods
A number of methods for sparse-matrix and operator recovery based on graph colorings have been proposed [curtis_powell_reid_74, coleman_more_83, stathopoulos_laeuchli_orginos_13, schafer_owhadi_21, frommer_schimmel_schweitzer_21, etc.]. [^4] These methods are sometimes called probing methods in the matrix function trace estimation literature. To the best of our knowledge, such methods were first considered in the 1970s, and are based on the following observation: Partition column indices $[d]$ into sets $\{C_{1},\ldots,C_{k}\}$ such that the columns of $\mathbf{A}$ corresponding to any given $C_{i}$ have disjoint support. Then we can recover all of the columns in a given set $C_{i}$ with a single matrix-vector product: a vector which is supported only on $C_{i}$. The sets $C_{i}$ can be obtained by coloring a graph. In particular, form a graph on $d$ vertices, where there is an edge between vertices $i$ and $j$ if and only if the $i$-th and $j$-th columns of $\mathbf{A}$ have overlapping support. A $k$-coloring of this graph gives the partition $\{C_{1},\ldots,C_{k}\}$.
@@ 4.1 Analysis of coloring methods on banded matrices
For any integer $b\geq 0$, let $\mathbf{S}\in\{0,1\}^{d\times d}$ be the sparsity pattern of banded matrices of bandwidth $s=2b+1$; that is, $[\mathbf{S}]_{i,j}=\mathds{1}(|i-j|\leq b)$. For convenience, we will assume $d=ks$, for some integer $k\geq 1$. This sparsity pattern yields a natural coloring-based partitioning of $[d]$ into the sets $C_{i}=\{i+js:j\in[k]\}$ for $i\in[s]$.
@@
Matrix-vector products with $\mathbf{A}^{(i)}$ can be computed with a single product to $\mathbf{A}$. Thus, using $s$ matrix-vector products, we obtain an unbiased estimator for $\mathbf{S}\circ\mathbf{A}$ with expected squared error $\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}$. Averaging $t$ independent copies of this estimator will reduce the variance by a factor of $t$. Hence, using $m\geq s/\varepsilon$ matrix-vector products, one obtains an algorithm with expected squared error bounded by $\varepsilon\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}$. While this algorithm is well-known in the literature [stathopoulos_laeuchli_orginos_13, frommer_schimmel_schweitzer_21, etc.], to the best of our knowledge, an analysis like the one described here has not been written down. As we discuss in the next section, this is marginally better than Theorem 1.
@@ 5 Numerical Experiments
In this section we provide several numerical experiments which illustrate the performance of Algorithm 1. These problems are modeled after similar problems from the literature. Code to reproduce the figures can be found at https://github.com/tchen-research/fixed_sparsity_matrix_approximation.
@@ 5.1 Model problem
We consider the matrix $\mathbf{A}=\mathbf{M}^{-1}$, where $\mathbf{M}=\operatorname{tridiag}(-1,4,-1)$. This class of matrices exhibits exponential decay away from the diagonal and was used in experiments in past work [benzi_simoncini_15, frommer_schimmel_schweitzer_21].
@@ 5.2 Trefethen Primes
We let $\mathbf{A}=\mathbf{M}^{-1}$, where $\mathbf{M}$ be the $1000\times 1000$ matrix whose entries are zero everywhere except for the primes $2,3,5,7,\ldots,7919$ along the main diagonal and the number 1 in all the positions $[\mathbf{B}]_{i,j}$ with $|i–j|\in\{1,2,4,8,\ldots,512\}$. [^5] In problem 7 of the “A Hundred-dollar, Hundred-digit Challenge” in SIAM News, readers are asked to compute the (1,1) entry of a larger but analogously defined matrix $\mathbf{A}$ to 100 digits of accuracy [trefethen_02]. An example (somewhat different from our example) involving this matrix was used in [park_nakatsukasa_23].
Refer to captionRefer to captionRefer to caption
@@
For $b>0$, we defined a sparsity pattern $\mathbf{S}$ to be such that, for each $t\in\{1,2,4,8,\ldots,512\}$, $[\mathbf{S}]_{i,j}=1$ whenever $|i-j\pm t|\leq b$ and zero otherwise. In other words, the sparsity pattern consists of bandwidth $12b+$ 1 bands centered nonzero entries of $\mathbf{M}=\mathbf{A}^{-1}$. This is illustrated in Figure 3.
The results are shown in Figure 4. Here, the convergence of Algorithm 1 is somewhat better than the upper bound Theorem 1. This is because many rows of the sparsity pattern have far fewer than $s$ entries. From the proof of Theorem 1, it is clear how to obtain an exact characterization of the expected squared error.
@@ 6 Outlook
It is clear that there is potential for combining coloring methods with algorithms such as Algorithm 1 in order to avoid paying for large off-sparsity entries in parts of a matrix while simultaneously maintaining the robustness to noise enjoyed by Algorithm 1. One approach to combining these two paradigms is to use adaptive queries to identify portions of the matrix with large-mass, and then find a coloring based on this information. We believe further study in this direction may yield algorithms which work better in many practical situations. Of course, our lower bound Theorem 2 shows that only constant factors can be improved for some families of problem instances.
The present paper focuses only on the Frobenius norm. It would be valuable to understand the analogous problem in other norms such as the matrix 2-norm. For other norms, $\mathbf{S}\circ\mathbf{A}$ is not necessarily the best approximation to $\mathbf{A}$ with sparsity $\mathbf{S}$. For instance, if $\mathbf{A}=[1,1;1,1]$ and $\mathbf{S}=[1,0;0,0]$, then
@@
Loosely speaking, minimizing the operator-norm approximation error requires the rows of the error matrix to be small and unaligned, whereas the Frobenius norm problem only requires them to be small.
@@ Acknowledgements
Appendix A Discussion on relative error recovery
Instead of Problem 2, one may hope for a guarantee like
$\|\mathbf{S}\circ\mathbf{A}-\widetilde{\mathbf{A}}\|_{\mathsf{F}}\leq\varepsilon\|\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}.$ (10)
Unfortunately, we cannot expect to be able to obtain such an approximation using fewer than $d$ queries in general; for $\varepsilon<1$, this would require differentiating matrices for which $\mathbf{S}\circ\mathbf{A}$ is zero from those for which it is arbitrarily small.
Even so, in many situations where one might hope to recover the sparse part of a matrix, $\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}$ can be bounded by a constant multiple of $\|\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}$. In such cases, Theorem 1 implies that Algorithm 1 returns an approximation satisfying Equation 10 using $O(s/\varepsilon^{2})$ queries. One instance where this can be expected is if $\mathbf{A}=f(\mathbf{M})$, where $\mathbf{M}$ has sparsity $\mathbf{S}$. It is known that matrix functions of sparse matrices have entries which decay exponentially away from the sparsity pattern [benzi_razouk_08]. Another instance in which this holds is if $\mathbf{S}$ contains the identity and $\mathbf{A}$ is diagonally dominant; i.e. if $|[\mathbf{A}]_{i,i}|\geq\sum_{j\neq i}|[\mathbf{A}]_{i,j}|$. Indeed, in this case
$\|\mathbf{A}-\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\leq\|\mathbf{A}-\mathbf{I}\circ\mathbf{A}\|_{\mathsf{F}}^{2}=\sum_{i=1}^{d}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{d}[\mathbf{A}]_{i,j}^{2}\leq\sum_{i=1}^{d}[\mathbf{A}]_{i,i}^{2}=\|\mathbf{I}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\leq\|\mathbf{S}\circ\mathbf{A}\|_{\mathsf{F}}^{2}.$
Using Lemma 3, we can derive lower-bounds. For instance, if in the proof of Theorem 2 we replace $F$ by
$F=\Big{\{}\|\mathbf{I}\circ\mathbf{A}\|_{\mathsf{F}}^{2}\leq b_{1}d^{3}\Big{\}}$
and $b_{1}$ by $150$, then by Fact 2, the proof of Theorem 2 can be repeated and results in a bound $\Omega(1/\varepsilon^{2})$ queries for relative error diagonal approximation of PSD matrices.
In fact, if we set $r$ as large constant multiple of $d$, then $\mathbf{A}$ becomes diagonally dominant with high probability, and the same lower bound still holds. In light of the above upper bound, Algorithm 1 is optimal for relative error diagonal approximation on diagonally dominant matrices.
Appendix B Lower Bound Lemmas
@@
B.1 Adaptive queries to Wishart matrices
In this section we provide a proof of Proposition 3, which is essentially [braverman_hazan_simchowitz_woodworth_20, Lemma 3.4]. This is mostly included for completeness. First, we consider what happens after a single non-adaptive query.
@@ B.1 Adaptive queries to Wishart matrices
Suppose $\mathbf{G}\sim\operatorname{Gaussian}(d,r)$. Let $\mathbf{x}$ be a unit-length query chosen independently of $\mathbf{G}$, and define $\mathbf{y}=\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{x}$. Then, there is an $n\times n$ orthonormal matrix $\hat{\mathbf{X}}=[\mathbf{x}~{}\mathbf{X}]$, constructed solely as a function of $\mathbf{x}$, and a matrix $\mathbf{H}\sim\operatorname{Gaussian}(d-1,r-1)$ independent of $\mathbf{x}$ and $\mathbf{y}$ such that such that
@@
The first row and column of this matrix depend only on $\mathbf{x}$ and $\mathbf{y}$. We will now show $\mathbf{X}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{G}\mathbf{X}$ is the sum of a matrix depending on $\mathbf{x}$ and $\mathbf{y}$ and Gaussian matrix independent of $\mathbf{x}$ and $\mathbf{y}$.
@@
$\hat{\mathbf{X}}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{G}\hat{\mathbf{X}}=\begin{bmatrix}\mathbf{x}^{\mathsf{T}}\mathbf{y}&\mathbf{y}^{\mathsf{T}}\mathbf{X}\\ \mathbf{X}^{\mathsf{T}}\mathbf{y}&(\mathbf{x}^{\mathsf{T}}\mathbf{y})^{-1}\mathbf{X}^{\mathsf{T}}\mathbf{y}\mathbf{y}^{\mathsf{T}}\mathbf{X}\end{bmatrix}+\begin{bmatrix}0&\mathbf{0}_{1,n-1}\\ \mathbf{0}_{n-1,1}&\mathbf{X}^{\mathsf{T}}\mathbf{G}^{\mathsf{T}}\mathbf{R}\mathbf{R}^{\mathsf{T}}\mathbf{G}\mathbf{X}\end{bmatrix}.$
It remains to show $\mathbf{R}^{\mathsf{T}}\mathbf{G}\mathbf{X}$ is a $(r-1)\times(d-1)$ Gaussian matrix independent of $\mathbf{x}$ and $\mathbf{y}$.
First, note that since ${\mathbf{X}}$ is chosen only based on $\mathbf{x}$ (which is independent of $\mathbf{G}$), $\mathbf{G}\hat{\mathbf{X}}$ consists of iid Gaussians. Thus, the columns of $\mathbf{G}{\mathbf{X}}$ are mutually independent of one another and $\mathbf{x}$, and hence $\mathbf{G}\mathbf{X}$ is mutually independent of $\mathbf{x}$ and $\mathbf{y}$. Finally, since $\mathbf{R}$ depends only on $\mathbf{r}=(\mathbf{x}^{\mathsf{T}}\mathbf{y})^{-1/2}\mathbf{G}\mathbf{x}$, $\mathbf{R}$ is independent of $\mathbf{G}\mathbf{X}$. Thus, $\mathbf{R}^{\mathsf{T}}\mathbf{G}\mathbf{X}$ has iid Gaussian entries independent of $\mathbf{x}$ and $\mathbf{y}$ (and hence any matrices constructed solely from $\mathbf{x}$ and $\mathbf{y}$). ∎
@@
We proceed by induction. Suppose that after $t$ queries, the result of the lemma holds. Let $\mathbf{x}_{t+1}$ be a query chosen based solely on $\mathbf{x}_{1},\ldots,\mathbf{x}_{t}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t}$ and hence independent of $\mathbf{G}_{t}$.
Then by the inductive hypothesis,
@@
Let $\mathbf{x}$ denote the bottom $d-t$ entries of $\mathbf{V}_{t}^{\mathsf{T}}\mathbf{x}_{t+1}$, normalized to have length 1. By Lemma 4, querying $\mathbf{G}_{t}^{\mathsf{T}}\mathbf{G}_{t}$ results in the factorization
@@
where $\mathbf{\Delta}$ and $\hat{\mathbf{X}}$ are constructed solely as functions of $\mathbf{x}$ and $\mathbf{G}_{t}^{\mathsf{T}}\mathbf{G}_{t}\mathbf{x}$ (and hence of $\mathbf{x}_{1},\ldots,\mathbf{x}_{t+1}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t+1}$), $\hat{\mathbf{X}}$ is orthonormal, and $\mathbf{H}\sim\operatorname{Gaussian}(d-t-1,r-t-1)$ is independent of $\mathbf{x}$ (and hence of $\mathbf{x}_{1},\ldots,\mathbf{x}_{t+1}$ and $\mathbf{y}_{1},\ldots,\mathbf{y}_{t+1}$).
@@
B.2 Anti-concentration for sums